what is the slope intercept form for 5x-9y=11
Found 2 solutions by mananth, TutorDelphia:
Answer by mananth(16946)

5x-9y=11
re write in slope intercept form
9y=5x-11
/9
y=(5/9)x-(11/9)
compare with y=mx+b
m= slope = (5/9)
y intercept = b = 11/9

Answer by TutorDelphia(193)

5x-9y=11
subtract 5x from both sides
-9y=-5x+11
divide both sides by -9
y=(5/9)x-11/9